OpenStudy (anonymous):

The first angle of a triangle is 10 degrees smaller than the second angle. The third angle is 10 degrees less than twice the second angle. Show that this is a right triangle.

OpenStudy (anonymous):

think of the angles as A B and C.

OpenStudy (anonymous):

A=b-10

OpenStudy (anonymous):

c=2b-10

OpenStudy (anonymous):

A+B+C=180

OpenStudy (anonymous):

that should help you solve it.

OpenStudy (anonymous):

Yeah, I get it know :) Thank you. I was taking the wrong approach before.

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

I set it up as b-10+2b-10=180 3b-20=180 3b=200 b=66.67?

OpenStudy (anonymous):

B=180-A-C B=180-(b-10)-(2b-10) B=180-3b-20 B=160-3b?

OpenStudy (math&ing001):

You forgot the third angle before : (b-10)+b+(2b-10)=180

OpenStudy (anonymous):

Oh.. 4b-20=180 4b=200 b=50?

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

then I plug it in and I get my other two angles right?

OpenStudy (math&ing001):

Exactly :)

OpenStudy (anonymous):

I got it, thank you :)

OpenStudy (math&ing001):

Welcome !